OpenStudy (anonymous):

evaluate : 4 − 14 ÷ 7 + 32

OpenStudy (ryan123345):

What do you think you do first?

OpenStudy (anonymous):

not sure

OpenStudy (ryan123345):

1) Parenthesis 2) Exponents 3) Multiplication / Division ( whichever comes first ) 4) Addition / Subtraction ( whichever comes first ) Welcome to Open Study by the way :)

OpenStudy (ryan123345):

lets look at number 1. Do we have any Parenthesis?

OpenStudy (anonymous):

division first

OpenStudy (anonymous):

is the amswer 30?

OpenStudy (ryan123345):

No but you got the first part right. 14 divided by 7 is 2 so 4 - 2 + 32

OpenStudy (ryan123345):

now in that equation ( 4 - 2 + 32 ) which comes first? Addition or Subtraction?

OpenStudy (anonymous):

actaully what i need help on is- Simplify the following expression. 4x − 6y + 2z − 5x − 3y + 8z

OpenStudy (ryan123345):

ill help you with that after. First tell me the answer of the first one.

OpenStudy (anonymous):

well i typed it wrong in the first place

OpenStudy (anonymous):

and understand it

OpenStudy (anonymous):

Simplify the following expression. 4x − 6y + 2z − 5x − 3y + 8z

OpenStudy (ryan123345):



OpenStudy (ryan123345):

answer those 3 questions I posted

OpenStudy (anonymous):

but how do you combine the like terms?

OpenStudy (anonymous):

never mind i see what you did ok hang on

OpenStudy (ryan123345):

you add letters with letters. you cant add something like 3z + 4y because they dont have the same letters

OpenStudy (ryan123345):

but you can add 3z + 3z because the variables are the same. It will just give you 6z

OpenStudy (anonymous):

-5x + 4x =-1x -6y + -3y =-9y 2z + 8z =10z

OpenStudy (ryan123345):

great job! so now you just put it out as an equation so you would get -x + -9y + 10z

OpenStudy (anonymous):

Thank you very much